14. The health care system wants to encourage a healthy lifestyle. All doctors are asked to speak with all their patients about healthy eating during every visit. Every week, a health care system takes a random sample of 100 patients seen in the week and asks them if the doctor they saw this week had brought up the topic of healthy eating. Across all weeks, the average percentage of visits in which healthy eating was brought up is 91 percent. What would be the upper control limit in the resulting p-chart? a. 0.0819 b. 0.004145 c. 0.175855 d. 0.9 e. None of the abovepls show calculation by hand.
Answer #1

The average visits in which the healthy eating was brought up= 0.91= Pm; n= 100

The average visits in which the healthy eating was brought up:0.91-Pm: n= 100 Upper control limit-Pm+3(1-Pm UCL= 0.91 + 3 /031(009) 7n 100 UCL-0.995

UCL= 0.995